# CheMFi: A Multifidelity Dataset of Quantum Chemical Properties of Diverse Molecules

### v1

@@ Abstract
Progress in both Machine Learning (ML) and conventional Quantum Chemistry (QC) computational methods have resulted in high accuracy ML models for QC properties ranging from atomization energies to excitation energies. Various datasets such as MD17, MD22, and WS22, which consist of properties calculated at some level of QC method, or fidelity, have been generated to benchmark such ML models. The term fidelity refers to the accuracy of the chosen QC method to the actual real value of the property. The higher the fidelity, the more accurate the calculated property, albeit at a higher computational cost.
@@
Research in multifidelity ML (MFML) methods, where ML models are trained on data from more than one numerical QC method, has shown the effectiveness of such models over single fidelity methods. Much research is progressing in this direction for diverse applications ranging from energy band gaps to excitation energies. A major hurdle for effective research in this field of research in the community is the lack of a diverse multifidelity dataset for benchmarking.
Here, we present a comprehensive multifidelity dataset drawn from the WS22 molecular conformations. We provide the quantum Chemistry MultiFidelity (CheMFi) dataset consisting of five fidelities calculated with the TD-DFT formalism. The fidelities differ in their basis set choice and are namely: STO-3G, 3-21G, 6-31G, def2-SVP, and def2-TZVP. CheMFi offers to the community a variety of QC properties including vertical excitation energies, oscillator strengths, molecular dipole moments, and ground state energies. In addition to the dataset, multifidelity benchmarks are set with state-of-the-art MFML and optimized-MFML.
Keywords: machine learning, multifidelity machine learning, dataset, Wigner sampling, quantum chemistry, DFT, orca software
1 Background and Summary
@@
Recent developments in the field of machine learning (ML) for quantum chemistry (QC) have significantly changed the landscape of research and discovery in QC properties 1, 2, 3, 4. This has allowed significant reduction in the time to calculate (or rather, predict) QC properties once an ML model has been trained. For such models, the protocol involves testing them against some benchmark datasets such as the MD17 5, QM7 6, or the QM9 dataset 7, 8. The MD17 dataset consists of trajectories of small to medium-sized molecules with their corresponding geometries, forces, and potential energies. In contrast, the QM7 and QM9 datasets are collections of various molecules and their isomers with properties ranging from dipole moments to enthalpies. These diverse datasets allow for a uniform assessment of ML methods for QC. Recently, the WS22 database was released with a collection of Wigner Sampled geometries of 10 diverse molecules 9, 10. With varied chemical complexity and number of atoms, the WS22 datasets provides a collection of QC properties for these molecules calculated at one level of theory, or fidelity. It was also shown that for this collection of molecules the use of ML methods is indeed challenging. This is primarily due to the wider chemical space that the Wigner Sampled geometries cover 9, 11.
@@
The usual method of training a ML model on a required level of QC theory has since been superseded by methodological improvements which harness inherent QC hierarchies to cancel out errors in different numerical QC methods. These are categorized as multifidelity methods. The most elementary multifidelity method, called $\Delta$-ML, involves the use of two QC methods, or fidelities, one significantly cheaper than the other 12. An ML model is trained on the difference between these two fidelities. Recently, a systematic generalization of the $\Delta$-ML method was implemented for excitation energies and termed multifidelity machine learning (MFML) 13 where more than two fidelities are used to train various ML models resulting in low-cost high-accuracy MFML model. Yet another multifidelity method which has been shown to be efficient in potential energy surface (PES) reconstruction is the hierarchical-ML (h-ML) approach 14. In h-ML, an ad hoc optimization is carried out to select the ideal number of training samples across various fidelities to build one composite multifidelity model. The optimized MFML method was introduced in Ref. 15 where the combination of the different sub-models built on different fidelities is optimized on a holdout validation set of reference properties. Such an o-MFML was shown to improve the accuracy of prediction for atomization energies and excitation energies across diverse molecules. Certain other flavors of ML using multifidelity data have been proposed and tested, including multi-task Gaussian processes treating the different fidelities as interdependent tasks 16, 17. Also, various multifidelity methods have been applied to predicting diverse QC properties such as band gaps in solids, excitation energies, and atomization energies of various molecules 18, 19, 20, 13.
@@
With significant push in the direction of research and development in the multifidelity methods, it becomes important to also follow protocols similar to the single fidelity ML methods. That is, there is a need for uniform benchmarking of multifidelity models for their accuracy of prediction, ease of transferability of the models, and reducing cost of generating training data among other things. Therefore, a diverse set of multifidelity data is needed to uniformly assess the various developed and developing multifidelity models in the area of ML for QC. One such dataset that exists is the QM7b 6, 21 which is an extension of the QM7 dataset. Originally created for multi-task learning, it has since been used in the methodological benchmarking of multifidelity models with atomization energies being calculated at three levels of theory with three varying basis set sizes 20, 15. Although the highest level of theory used in this dataset is the gold standard Coupled Cluster Singles and Doubles perturbative Triples (CCSD(T)), it provides a limited possibility to assess the effectiveness of multifidelity methods due to its small size, 7,211 samples. This is indeed a small amount of data to enable proper benchmarking of multifidelity methods. In addition to this, the QM7b is a collection of only small to medium sized molecule with the number of atoms ranging from 4 to 23. Further, as recent studies have shown, multifidelity methods tend to behave differently for different properties ranging from atomization energies to excitation energies to band gaps in polymers 20, 19, 15. This results in independent and uncorrelated developments of the multifidelity method which although displaying benefit for one property might fall short, while being used for another QC property.
@@
To unify the research in this rapidly developing field of multifidelity methods, it becomes necessary to present to the community a diverse collection of multifidelity data over a range of molecular complexity. Building up on existing datasets is preferred in such a scenario to prevent redundant calculations and geometry generation. After all, the entire point of a multifidelity method is to reduce compute cost and resource usage in discovery and research. In interest of such an approach, the WS22 database 10 was chosen to be the collection of geometries. In addition to being a collection of molecules that are chemically complex with distinct conformers, the molecule in this dataset also cover a wide range of the quantum chemical configuration space in contrast to other datasets such as MD17. The presence of flexible functional groups make the geometries, and by extension, the QC properties, of this dataset challenging for ML models to learn. These features make this collection the preferred choice to generate multifidelity data. For each of the molecules of increasing size and chemical complexity, this dataset offers 120,000 geometries. This creates a vast dataset collection of diverse geometries covering various conformers of the different molecules. In total there are around 1 million geometries in the WS22 database. Performing multifidelity QC calculations for such a vast number of geometries is not feasible. It is more realistic and computationally feasible to produce a multifidelity dataset for a portion of the geometries of the WS22 database. Therefore, for each of the molecules in the WS22 database, 15,000 geometries were evenly sampled (see supplementary material section S1 for more details) and the multifidelity QC calculations performed for these. Thus, in total, $9\times 15,000=135,000$ geometries were sampled and the multifidelity data was computed for these. This dataset is provided to the ML-QC community under the name CheMFi (quantum Chemistry MultiFidelity) dataset 22. A detailed description of the geometry sampling, data generation procedure, the fidelities, and the technical details of the CheMFi dataset are provided in the following section. In addition, scripts to generate two multifidelity models from Ref. 15 are provided. These are multifidelity machine learning (MFML) and optimized MFML (o-MFML) models. The ML methods such as MFML and o-MFML are discussed in the supplementary materials in Section S2 to retain readability of this main text.
@@
The diverse collection of molecules in CheMFi along with their multifidelity properties, provides a challenging dataset for the domain of ML in QC. Due to the large number of multifidelity data points, and easily usable associated scripts, we believe that CheMFi is a significant collection that will help push the boundaries of multifidelity methods for ML in QC properties.
2 Methods
Figure 1: The workflow of generating the CheMFi dataset by sampling from the WS22 database. 15,000 geometries are used for each molecule resulting in a total of 135,000 single point geometries. For each of these, multiple QC properties are calculated at DFT level of theory with varying basis set sizes to create the diverse multifidelity dataset. Molecular geometry images taken from Wikimedia Commons and Ref. 9.
@@ 2 Methods
In addition to these molecules, toluene is also included to compare with the MD17 5 database. Since toluene consists of a single conformer and was only introduced in WS22 for comparison to existing datasets such as MD17, this molecule was not included while generating the CheMFi dataset. The original WS22 database was first generated as reported extensively in Ref. 9. The pipeline involves optimized equilibrium geometries identification for the different conformations of the molecule with DFT 23, 24. Following this, the respective Wigner Sampling is carried out from ground state (S0) and/or excited state (S1) minima. For these, the geometries are subsequently interpolated by finding on a Riemann manifold, an optimized geodesic curve. The metric for this is defined by a redundant internal coordinate functions 25. In the original WS22 database provided in Ref. 10, this results in a little over 1 million samples across 9 molecules with various properties calculated at the TD-DFT level of theory using the PBE0/6-311 G* functional and basis set combination 9.

### v2

@@ Abstract
Progress in both Machine Learning (ML) and Quantum Chemistry (QC) methods have resulted in high accuracy ML models for QC properties. Datasets such as MD17 and WS22 have been used to benchmark these models at some level of QC method, or fidelity, which refers to the accuracy of the chosen QC method. Multifidelity ML (MFML) methods, where models are trained on data from more than one fidelity, have shown to be effective over single fidelity methods. Much research is progressing in this direction for diverse applications ranging from energy band gaps to excitation energies. One hurdle for effective research here is the lack of a diverse multifidelity dataset for benchmarking. We provide the quantum Chemistry MultiFidelity (CheMFi) dataset consisting of five fidelities calculated with the TD-DFT formalism. The fidelities differ in their basis set choice: STO-3G, 3-21G, 6-31G, def2-SVP, and def2-TZVP. CheMFi offers to the community a variety of QC properties such as vertical excitation properties and molecular dipole moments, further including QC computation times allowing for a time benefit benchmark of multifidelity models for ML-QC.
Background & Summary
@@
Recent developments in the field of machine learning (ML) for quantum chemistry (QC) have significantly changed the landscape of research and discovery in QC properties [1, 2, 3, 4] with significant reduction in the time to predict QC properties once an ML model has been trained. For such models, the protocol often involves testing them against some benchmark datasets such as the MD17 [5], QM7 [6], or the QM9 dataset [7, 8]. Recently, the WS22 database was released with a collection of Wigner Sampled geometries of 10 diverse molecules [9, 10]. With varied chemical complexity and number of atoms, the WS22 datasets provides a collection of QC properties for these molecules calculated at one level of theory, or fidelity. It was also shown that for this collection of molecules the use of ML methods is indeed challenging due to the wider chemical space that the geometries cover [9, 11].
@@
Multifidelity methods harnessing inherent QC hierarchies to cancel out errors across different numerical QC methods have since superseded the single fidelity ML methods. These methods include $\Delta$-ML [12] based models such as hierarchical machine learning [13], multifidelity machine learning (MFML) [14, 15], and optimized MFML (o-MFML) [16]. Certain other flavors of ML using multifidelity data have been proposed and tested, including multi-task Gaussian processes treating the different fidelities as interdependent tasks [17, 18]. Multifidelity methods have been applied to predicting diverse QC properties such as band gaps in solids, excitation energies, and atomization energies of various molecules [19, 20, 14, 15].
@@
With growing interest in multifidelity methods, there is a need to establish uniform benchmarking protocols to evaluate prediction accuracy, transferability, and training data cost-efficiency, similar to the single fidelity ML methods. Therefore, a diverse set of multifidelity data is needed to uniformly assess the various developed and developing multifidelity models in the area of ML for QC. One such dataset that exists is the QM7b [6, 21] which is an extension of the QM7 dataset. Originally created for multi-task learning, it has since been used in the methodological benchmarking of multifidelity models with atomization energies being calculated at three levels of theory with three varying basis set sizes [14, 16]. Although the highest level of theory used in this dataset is the gold standard Coupled Cluster Singles and Doubles perturbative Triples (CCSD(T)), it provides a limited possibility to assess the effectiveness of multifidelity methods due to its small size of 7,211 samples. The lack of QC-computation times for the different fidelities, prevent a time-benefit assessment of multifidelity methods on this dataset. In addition, the QM7b is a collection of only small to medium sized molecule with the number of atoms ranging from 4 to 23.
@@
To unify the research in this rapidly developing field of multifidelity methods, it becomes necessary to present to the community a diverse collection of multifidelity data over a range of molecular complexity. Building up on existing datasets is preferred in such a scenario to prevent redundant calculations and geometry generation. After all, the entire point of a multifidelity method is to reduce compute cost and resource usage in discovery and research. In interest of such an approach, the WS22 database [10] was chosen to be the collection of geometries. In addition to being a collection of molecules that are chemically complex with distinct conformers, the molecule in this dataset also cover a wide range of the quantum chemical configuration space in contrast to other datasets such as MD17. The presence of flexible functional groups make the geometries, and by extension, the QC properties, of this dataset challenging for ML models to learn. These features make this collection the preferred choice to generate multifidelity data. For each of the molecules of increasing size and chemical complexity, this dataset offers 120,000 geometries. This creates a vast dataset collection of diverse geometries covering various conformers of the different molecules. In total there are around 1 million geometries in the WS22 database. Performing multifidelity QC calculations for such a vast number of geometries is not feasible. It is more realistic and computationally feasible to produce a multifidelity dataset for a portion of the geometries of the WS22 database. Therefore, for each of the molecules in the WS22 database, 15,000 geometries were evenly sampled, for a total of $9\times 15,000=135,000$ geometries, and the multifidelity QC calculations performed for these.
@@
This dataset is provided to the ML-QC community under the name CheMFi (quantum Chemistry MultiFidelity) dataset [22]. A detailed description of the geometry sampling, data generation procedure, the fidelities, and the technical details of the CheMFi dataset are provided in the following section. In addition, scripts to generate two multifidelity models from ref. [16], namely, MFML and optimized MFML (o-MFML) are provided. These ML methods are discussed in the Methods section of this work. Scripts to assess time benefit of multifidelity methods are also included. This makes it easy for future research in the multifidelity methods to establish a clear time benefit for these models over standard single fidelity ML methods. The diverse collection of molecules in CheMFi along with their multifidelity properties, provides a challenging dataset for the domain of ML in QC. Due to the large number of multifidelity data points, and easily usable associated scripts, we believe that CheMFi is a significant collection that will help push the boundaries of multifidelity methods for ML in QC properties.
Methods
@@
In addition to these molecules, toluene is also included to compare with the MD17 [5] database. Since toluene consists of a single conformer and was only introduced in WS22 for comparison to existing datasets such as MD17, this molecule was not included while generating the CheMFi dataset. The original WS22 database was first generated as reported extensively in ref. [9]. The pipeline involves optimized equilibrium geometries identification for the different conformations of the molecule with DFT [23, 24]. Following this, the respective Wigner Sampling is carried out from ground state (S0) and/or excited state (S1) minima. For these, the geometries are subsequently interpolated by finding on a Riemann manifold, an optimized geodesic curve. The metric for this is defined by a redundant internal coordinate functions [25]. In the original WS22 database provided in ref. [10], this results in a little over 1 million samples across 9 molecules with various properties calculated at the TD-DFT level of theory using the PBE0/6-311 G* functional and basis set combination [9].
Data Sampling and Quantum Chemistry Calculations